# Discovery and Expansion of New Domains within Diffusion Models

### v2

@@ 4.1 Experimental Setup
Model Zoos and Datasets. We adopt four pre-trained DDPMs on different single domain datasets as our base models for experiments: improved DDPM [36] trained on AFHQ-Dog [8], and DDPM [19] trained on CelebA-HQ [26], LSUN-Church [65], and LSUN-Bedroom [65]. Each model generates images in the original resolution of $256\times 256$, resulting in a total dimensionality of the latent spaces $d=256\times 256\times 3=196,608$.
@@
In addition to the above commonly used natural image datasets, we further experiment with two astrophysical datasets to cover a wide range of domain differences and to showcase the application scenarios with scientific data. Specifically, we adopt the GalaxyZoo [58] and the radiation simulation data [62], the latter has been investigated using DMs for prediction purposes. Details about those astrophysical datasets, their scientific interpretations, and evaluations are included in the Appendix F for interested readers, which differs from the usual interpretation of natural images.
@@
Comparisons. Our main experiments focus on the comparison with the trajectory tuning-based approaches using the given DMs and OOD samples. Specially, we compare with vanilla tuning, where we finetune the pre-trained model with classic VLB loss [19] on unseen images, as well as several different SOTA image-to-image translation methods via diffusion models as baselines, as they also output images in domains that are different to the trained ones. However, it is worth noting that those baseline methods (i.e., DiffusionClip [29], and Asyrp [31]) are not really generating unseen images, but rather editing an given original ID image to a target unseen domain. Moreover, those are learning-based methods trained on each unseen domain with extra CLIP loss [39], while our method operates on a single mutual latent space from the pre-trained base diffusion models without additional external supervision, echoing our conceptual design of “space expansion”.
@@
Implementations and Learning Budget. We use $N=1000$ images for OOD domains. For deterministic diffusion, we adopt the standard DDIMs skipping step technique to accelerate both processes using 60 steps in total. Each direction takes an average of 3 - 6 seconds. We use the univariant Gaussian estimation for inverted OOD latent encodings. The geometric optimization tolerance $\omega$ is set to be 0.3 for distance for 0.1 for angles, which leads to a rejection rate of approximately 84.44 % based on initial samples. We use 2 RTX 3090 GPUs for all experiments including baselines. For baseline methods that perform image translation and editing [29, 31], we use their respective officially released implementations, each tuning process takes approximately 30 mins, with an initial default learning rate of 2e-6. To ensure a fair comparison, we also stop the vanilla tuning after around 30 minutes, resulting in around 20 epochs of finetuning, with an even larger learning rate of 1e-5. For our proposed method, the inversion takes the same time as those baseline methods, but the core proposed sampling methods take negligible time.
4.2 Results, Evaluations, and Analysis
Table 2: General quality evaluation in cross model and domain setup. T stands for Tuning-based, and TF denotes Tuning-Free. We report the FID scores ($\downarrow$) for natural image domains and the Mean Opinion Scores (MOS) ($\uparrow$) from subjective evaluations with astrophysicists. Note that most baseline methods perform the image-to-image translation and use additional CLIP loss to tune the model, and thus largely facilitate the task by bypassing the actual sampling stage and with extra strong semantic guidance. We therefore call for special attention when comparing the scores for a comprehensive and objective assessment, qualitative examples in Fig. 1 and more samples in Appendix F.
Methods Category Dog CelebA Church Bedroom Galaxy Radiation
Vanilla tuning T 213.6$\pm$4.8 229.7$\pm$4.3 192.5$\pm$3.7 191.1$\pm$4.0 - -
DiffusionCLIP [29] T 73.6$\pm$2.9 63.6$\pm$3.0 66.3$\pm$2.8 68.1$\pm$2.8 - -
Asyrp [31] T 66.4$\pm$2.1 58.7$\pm$2.4 62.5$\pm$2.0 63.2$\pm$1.9 1.84$\pm$0.97 1.35$\pm$0.74
Ours TF 78.2$\pm$2.8 64.5$\pm$2.7 64.8$\pm$2.7 62.9$\pm$2.6 2.88$\pm$0.93 1.52$\pm$0.80
@@ 4.2 Results, Evaluations, and Analysis
General Quality. As a general quality evaluation, we calculate the FID scores [18] on 5K generated samples for natural images and report the Mean Opinion Scores (MOS) for astrophysical data. The FID scores are averaged over four DDPMs pre-trained on different image domains, and the Mean Opinion Scores (MOS) with a scale between 1-5 are collected from subjective evaluations performed by astrophysicists with respect to the ground truth observation and simulation data in a non cherry-picky manner. As shown in Fig. 1 and Tab. 2, vanilla tuning with only image supervisions can hardly alter the original generation trajectories and synthesize desired images, always synthesizing in-domain images after comparable tuning time with other tuning baselines. As for methods that finetune the model with additional CLIP loss [39], such as DiffusionCLIP [29] and Asyrp [31], they relatively perform better for domains closer to their trained domains as expected. Our proposed method shows an opposite trend by achieving better performance in data domains with bigger differences, as it is easier to avoid mode interference with larger domain gaps in the latent spaces.
Data Diversity. The trade-off between the “general quality” and “diversity” of the synthesized data has long been a dilemma in the generative modeling study, as one can always trade one for another. For natural image generation, data diversity is a popular evaluation criteria; for scientific data synthesis, however, the embedded physical properties, rather than the perceptual difference, are usually more appreciated. As our proposed data sampling method in Sec. 3 deploys one inverted latent encoding $\mathbf{x}_{od,t}$ as a reference
Figure 4: Illustration of the data diversity between the latent reference and sampled data. Dogs belong to the ID case, while the remaining ones are OODs.
before exploring new samples, one may be concerned that the synthesized data will look similar to the reference image after denoising. To address the above concern, we qualitatively show examples of the reference and sampled data after denoising for both ID and OOD cases in Fig. 4. We observe that, when applied in an ID scenario, our proposed tuning-free method can also synthesize data with sufficient diversity. The diversity is relatively reduced in closer natural image domains, but it becomes easier to synthesize more diverse data with larger domain gaps, consistent with the trend observed in general quality assessments. We also note that for tuning baselines with extra CLIP supervisions [29, 31], the data diversity is not explicitly used as an evaluation criterion given their task objective of data editing. On the contrary, the data after editing (is expected to) preserve very high similarity in visual appearance, but with only difference in semantic attribute.
Impact of the Latent Step $t$. We test the latent step $t$ of diffusion models where we perform the latent inversion and OOD sampling as part of the ablation studies. As an overall takeaway, the choice of $t$ also characterizes a trade-off between the sampling difficulty and the risk of “mode interference”. In other words, when $t$ gets closer to the departure latent space $\mathcal{X}_{T}$, it is usually easier to conduct initial sampling as the inverted prior is closer to theoretical Gaussian but also tends to get more easily captured by the original trained domain as the ID latent spaces get larger probabilistic support in $\mathcal{X}_{T}$. Due to the space limit, we present other ablation studies that discuss the impact of the DDPM model designs and the degree of stochasticity in Appendix F.
5 Related Work
This work is closely related to several research fields such as the generalization ability of generative models [54, 43], the diffusion models and their deterministic variants [47, 49, 48, 51], the study on the latent space of deep generative models [26, 1, 12, 74], as well as recent downstream applications that use diffusion models for scientific explorations in a data-rare cases [62, 63]. Our work adopts a few-shot scenario to study the generalization abilities, uses the deterministic variant as the tool to achieve a bidirectional transition between latent noisy and data spaces, and contributes to a better understanding of those latent spaces. An extended discussion on related work is in Appendix B.
6 Further Discussions
@@
Conclusion. To sum up, we study the generalization abilities of DDPMs in the few-shot scenario. From the analytical point of view, we explore the generalization properties of DDPMs on unseen OOD domains. From the methodological perspective, our analytical results allow us to propose a novel paradigm for synthesizing images from new domains without tuning the generative trajectories. We also showcase the superiority of our method in data-sparse cases with large domain gaps.
Limitations and Future Directions. The current limitations and challenge mainly come from the OOD sampling. As previously discussed, the sampling from inverted OOD prior in high dimensionality is a challenging task and open research question, which also directly impact the synthesized image quality. Improved sampling methods are worth investigating as future research directions.
Broader Impact. This work falls into the category of generative models and their applications, we thus acknowledge that it may pose the same risks of malicious use of synthetic data as other general generative works. We impose necessary safeguards to ensure the responsible usage and distribution of data. However, the primary objective of our work is not performance-driven but to provide a better understanding of the generalization properties of diffusion generative models.
Appendix A Appendix / supplemental material
We structure the appendices as follows: We provide a detailed discussion on the related work in Appendix B. In Appendix C, we present the background of deterministic diffusion models. In Appendix D, we provide detailed proof for the lemmas used in Sec. 2 of the main paper as part of analytical studies. Meanwhile, Appendix D includes the necessary background on the geometric properties of high-dimensional Gaussians. More latent sampling methods are discussed in Appendix E. Appendix F includes additional details about the generative experiments on unseen OOD domains.
@@ Appendix B Related Work
Due to the space limitations in the main paper, we present a detailed discussion of related work in our appendices.
@@ B.1 Generalization in Generative Models
Domain Generalization [54] that aims to generalize models’ ability to extended data distributions has been an important research topic in broad machine learning area [13, 68, 70, 35, 33], with various computer vision applications such as recognition [37, 41], detection [66] and segmentation  [22, 14]. In the vision generative field, it becomes an even more challenging task with the extra demand to sample from the generalized distributions. One popular recent trend in the computer vision community is scaling up the model and dataset sizes as the most intuitive and obvious solutions [40, 20, 43]. Another scenario to study the domain generalization of generative models is within the few-shot scenario, where we only have a limited amount of data compared to the training set. In this case, fine-tuning the given model on the limited images [29] is the most straightforward way to go.
@@ B.2 Diffusion Models and Deterministic Variants
Diffusion Models (DMs) [47, 19, 49] are the state-of-the-art generative models for data synthesis in images [40, 43, 36, 16, 11, 23], videos [21], and audio [30, 75, 34]. There are currently two mainstream fundamental formulations of diffusion models, i.e., the denoising diffusion probabilistic models (DDPMs) [19] and score-based models [50]. One common perspective to understand both formulations is to consider the data generation as solving stochastic differential equations (SDEs), which characterize a stochastic process. Based on vanilla models, both branch develops their own deterministic variants, i.e., denoising diffusion implicit models (DDIMs) [48] and consistency models [51], with their core idea to follow the marginal distributions in denoising. Compared to initial DDPMs and Score-based DMs with ancestral sampling, the deterministic variants are solving ODEs instead of SDEs and largely accelerate the generation speed with fewer steps.
We leverage the deterministic variant (DDIMs [48]) as the tool to achieve bidirectional transition between latent noisy space and data space in this work.
@@ B.3 Latent Space of Deep Generative Models
Comprehensive studies of latent space of generative models [26, 1, 12] help to better understand the model and also benefit downstream tasks such as data editing and manipulation [73, 45, 31, 72]. A large portion of work has been exploring this problem within the context of GAN inversion [61], where the typical methods can be mainly divided into either learning-based [73, 42, 57, 3] or optimization-based categories [1, 2, 24, 10]. More recently, with the growing popularity of diffusion models, researchers have also focused on the latent space understanding of DMs for better synthesis qualities or semantic control [43, 74, 64].
@@ B.4 Diffusion Models in Science
While DMs have been extensively applied in data generation and editing within the multimodal context [43, 21, 75, 64, 74], recent works have extended their application domains to scientific explorations, such as astrophysics [63, 62], medical imaging [28, 59], and biology [60]. Compared to conventional computer vision applications, scientific tasks usually exhibit several distinct features. For instance, data acquisition and annotation are generally more expensive due to their scientific nature, resulting in a relatively smaller amount of available data for experiments. Additionally, the evaluation of these works adheres to established conventions within their respective contexts, which are usually different from image synthesis evaluation based on perceptual quality.
@@
Appendix C Deterministic Diffusion
Our analytical studies and methodology designs are built upon a specific variant of diffusion formulations, i.e., the deterministic diffusion process. While the original DDPMs involve a stochastic process for data generation via denoising (i.e., the same latent encoding will output different denoised images every time after the same generative chain), there is a variant of diffusion model that allows us to perform the denoising process in a deterministic way, known as the Denoising Diffusion Implicit Models (DDIMs) [48]. DDIMs were initially proposed for the purpose of speeding up the denoising process, however, later research works extend DDIMs from faster sampling application to other usages including the inversion technique to convert a raw image to its arbitrary latent space in a deterministic and tractable way. As briefly stated in our main paper, the core theoretical difference between DDIMs and DDPMs lies within the nature of forward process, which modifies a Markovian process to a non-Markovian one.
@@ Appendix C Deterministic Diffusion
The choice of mean function from Eqn. 6 ensures that $q_{\sigma}(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\sqrt{\alpha_{t}}\mathbf{x}_{0},(1-\alpha_{t})\mathbf{I})$ for all $t$, so that it defines a joint inference distribution that matches the “marginals” as desired. The non-Markovian forward process can be derived from Bayes’ rule:
@@
In fact, in the original paper, the authors also explicitly stated that: “ The forward process from Eqn. 7 is also Gaussian (although we do not use this fact for the remainder of this paper)”.  [^6] This paper refer to the DDIM paper [48]. While this Gaussian property was not emphasized and leveraged in the DDIMs paper, we find it useful in our context to explore the representation and generalization ability of pre-trained DDPMs.
In particular, the hyper-parameters for Gaussian scheduling $\alpha$ and $\beta$ in the context of DDIMs are slightly different from the original formulation in DDPMs [19]. Denote the original sequences from DDPMs as $\alpha^{\prime}_{t}$, then the $\alpha_{t}$ in this work follows the definition of DDIMs to be $\alpha_{t}=\prod^{T}_{t=1}\alpha^{\prime}_{t}$.
In addition to DDIMs, we note that the score-based formulation has also recently marked a deterministic variant, namely the Consistency Models [51]. The core idea of the consistency model is, to some extent, similar to DDIMs, which allows the vanilla score-based stochastic diffusion models to achieve “one-step” denoising, by following the marginal distributions.
@@
Table 3: Reconstruction results for arbitrary images via deterministic diffusion. We use an iDDPM [36] trained on AFHQ-Dog and 1K testing OOD images to compute the MAE (mean absolute error) reconstruction metric. Note DDIMs [48] was initially proposed to accelerate DDPMs sampling, but have not been studied in this OOD reconstruction setting.
@@
pSp [42] CelebA (ID) 0.079
e4e [52] CelebA (ID) 0.092
ReStyle [3] CelebA (ID) 0.089
HFGI [55] CelebA (ID) 0.062
DDIMs [48] Dog (ID) 0.073 $\pm$ 6e-4
@@
Appendix D Detailed Proofs and Geometric Properties
In this section, we provide detailed proof for the lemmas in Sec. 2. Particularly, Lemma 2.2 is a known property in high-dimensional Gaussian studies.
D.1 Proof of Lemma 2.1
We restate the lemma below, and provide the detailed proof, which has been introduced in the original DDIM paper [48].
Lemma D.1.

### v3

@@ 3.1 Experimental Setup
Model Zoos and Datasets. We adopt four pre-trained DDPMs on different single domain datasets as our base models for experiments: improved DDPM [nichol2021improveddmp] trained on AFHQ-Dog [choi2020stargan], and DDPM [ho2020dpm] trained on CelebA-HQ [karras2017progressive], LSUN-Church [yu2015lsun], and LSUN-Bedroom [yu2015lsun]. Each model generates images in the original resolution of $256\times 256$, resulting in a total dimensionality of the latent spaces $d=256\times 256\times 3=196,608$.
@@
In addition to the above commonly used natural image datasets, we further experiment with two astrophysical datasets to cover a wide range of domain differences and to showcase the application scenarios with scientific data. Specifically, we adopt the GalaxyZoo [willett2013galaxy] and the radiation simulation data [xu2023predicting], the latter has been investigated using DMs for prediction purposes. Details about those astrophysical datasets, their scientific interpretations, and evaluations are included in the Appendix E for interested readers, which differs from the usual interpretation of natural images.
@@
Comparisons. Our main experiments focus on the comparison with the trajectory tuning-based approaches using the given DMs and OOD samples. Specially, we compare with vanilla tuning, where we finetune the pre-trained model with classic VLB loss [ho2020dpm] on unseen images, as well as several different SOTA image-to-image translation methods via diffusion models as baselines, as they also output images in domains that are different to the trained ones. However, it is worth noting that those baseline methods (i.e., DiffusionClip [kim2022diffusionclip], and Asyrp [kwon2022diffusion]) are not really generating unseen images, but rather editing an given original ID image to a target unseen domain. Moreover, those are learning-based methods trained on each unseen domain with extra CLIP loss [radford2021clip], while our method operates on a single mutual latent space from the pre-trained base diffusion models without additional external supervision, echoing our conceptual design of “space expansion”.
@@
Implementations and Learning Budget. We use $N=1000$ images for OOD domains, and the univariant Gaussian estimation for inverted OOD latent encodings. We use 2 RTX 3090 GPUs for all experiments including baselines. For baseline methods that perform image translation and editing [kim2022diffusionclip, kwon2022diffusion], we use their respective officially released implementations, each tuning process takes approximately 30 mins, with an initial default learning rate of 2e-6. To ensure a fair comparison, we also stop the vanilla tuning after around 30 minutes, resulting in around 20 epochs of finetuning, with an even larger learning rate of 1e-5. For our proposed method, the inversion takes the same time as those baseline methods, but the core proposed sampling methods take negligible time.
@@
Main Results. As the general quality evaluation, we calculate the FID scores [heusel2017fid] for natural images and report the Mean Opinion Scores (MOS) for astrophysical data. The FID scores are averaged over four DDPMs pre-trained on different image domains, and the Mean Opinion Scores (MOS) with a scale between 1-5 are collected from subjective evaluations performed by astrophysicists with respect to the ground truth observation and simulation data in a non cherry-picky manner. As shown in Fig. 2 and Tab. 1, vanilla tuning with only image supervisions can hardly alter the original generation trajectories and synthesize desired images, always synthesizing in-domain images after comparable tuning time with other tuning baselines. As for methods that finetune the model with additional CLIP loss [radford2021clip], such as DiffusionCLIP [kim2022diffusionclip] and Asyrp [kwon2022diffusion], they relatively perform better for domains closer to their trained domains as expected. Our proposed method shows an opposite trend by achieving better performance in data domains with bigger differences, as it is easier to avoid mode interference with larger domain gaps in the latent spaces.
Table 1: General quality evaluation in cross model and domain setup. T stands for Tuning-based, and TF denotes Tuning-Free. We report the FID scores ($\downarrow$) for natural image domains and the Mean Opinion Scores (MOS) ($\uparrow$) from subjective evaluations with astrophysicists. Note that most baseline methods perform the image-to-image translation and use additional CLIP loss to tune the model, and thus largely facilitate the task by bypassing the actual sampling stage and with extra strong semantic guidance. We therefore call for special attention when comparing the scores for a comprehensive and objective assessment, qualitative examples in Fig. 2.
Methods Dog CelebA Church Bedroom Galaxy Radiation
[…]
4 Conclusion
@@
To sum up, we study the generalization abilities of diffusion models in the few-shot scenario. From the analytical point of view, we explore the generalization properties of diffusion models on unseen OOD domains. From the methodological perspective, our analytical study allows us to propose a sampling-based method for synthesizing images from new domains without tuning the pre-trained generative trajectories. In addition to experiments on natural images, we also showcase the superiority of our method in data-sparse cases with large domain gaps, such as in astrophysics.
Acknowledgments and Disclosure of Funding
This research was primarily conducted while YZ was a postdoctoral researcher at Princeton University. YZ also acknowledges travel funding from the French National Research Agency (ANR) via the “GraspGNNs” JCJC grant (ANR-24-CE23-3888), coordinated by Johannes F. Lutzeyer from École Polytechnique.
Appendix A Technical Appendices and Supplementary Material
We structure the appendices as follows: We provide a detailed discussion on the related work in Appendix B. In Appendix C, we present additional analysis of the latent OOD properties, including the background of deterministic diffusion models, the reconstruction test, the Gaussianity and the separability from the original ID latent prior.
@@ B.1 Generalization in Generative Models
Domain Generalization [wang2022generalizing] that aims to generalize models’ ability to extended data distributions has been an important research topic in broad machine learning area [ganin2016domain, zhao2020domain, zhou2021domain, muandet2013domain, li2017deeper], with various computer vision applications such as recognition [peng2019moment, rebuffi2017learning], detection [zhang2021understanding] and segmentation  [hoffman2018cycada, gong2019dlow]. In the vision generative field, it becomes an even more challenging task with the extra demand to sample from the generalized distributions. One popular recent trend in the computer vision community is scaling up the model and dataset sizes as the most intuitive and obvious solutions [ramesh2022dalle2, ho2022imagen, stablediff]. Another scenario to study the domain generalization of generative models is within the few-shot scenario, where we only have a limited amount of data compared to the training set. In this case, fine-tuning the given model on the limited images [kim2022diffusionclip] is the most straightforward way to go.
@@ B.2 Diffusion Models and Deterministic Variants
Diffusion Models (DMs) [sohl2015dpm_thermo, ho2020dpm, song2019generative] are the state-of-the-art generative models for data synthesis in images [ramesh2022dalle2, stablediff, nichol2021improveddmp, gu2021vector, dhariwal2021diff-img1, hu2021global], videos [ho2022video], and audio [kong2020diffwave, zhu2022discrete, mittal2021symbolic-diff]. There are currently two mainstream fundamental formulations of diffusion models, i.e., the denoising diffusion probabilistic models (DDPMs) [ho2020dpm] and score-based models [theo-diff2]. One common perspective to understand both formulations is to consider the data generation as solving stochastic differential equations (SDEs), which characterize a stochastic process. Based on vanilla models, both branch develops their own deterministic variants, i.e., denoising diffusion implicit models (DDIMs) [song2020ddim] and consistency models [song2023consistency], with their core idea to follow the marginal distributions in denoising. Compared to initial DDPMs and Score-based DMs with ancestral sampling, the deterministic variants are solving ODEs instead of SDEs and largely accelerate the generation speed with fewer steps.
We leverage the deterministic variant (DDIMs [song2020ddim]) as the tool to achieve bidirectional transition between latent noisy space and data space in this work.
@@ B.3 Latent Space of Deep Generative Models
Comprehensive studies of latent space of generative models [karras2017progressive, abdal2019image2stylegan, gal2022stylegan] help to better understand the model and also benefit downstream tasks such as data editing and manipulation [zhu2016generative, shen2020interpreting, kwon2022diffusion, zhu2020domain]. A large portion of work has been exploring this problem within the context of GAN inversion [xia2022gan], where the typical methods can be mainly divided into either learning-based [zhu2016generative, richardson2021encoding, wei2022e2style, alaluf2021restyle] or optimization-based categories [abdal2019image2stylegan, abdal2020image2stylegan++, huh2020transforming, creswell2018inverting]. More recently, with the growing popularity of diffusion models, researchers have also focused on the latent space understanding of DMs for better synthesis qualities or semantic control [stablediff, zhu2023boundary, yang2023cow].
@@ B.4 Diffusion Models in Science
While DMs have been extensively applied in data generation and editing within the multimodal context [stablediff, ho2022video, zhu2022discrete, yang2023cow, zhu2023boundary], recent works have extended their application domains to scientific explorations, such as astrophysics [xu2023denoising, xu2023predicting], medical imaging [kazerouni2023diffusion, wu2024medsegdiff], and biology [wu2209protein]. Compared to conventional computer vision applications, scientific tasks usually exhibit several distinct features. For instance, data acquisition and annotation are generally more expensive due to their scientific nature, resulting in a relatively smaller amount of available data for experiments. Additionally, the evaluation of these works adheres to established conventions within their respective contexts, which are usually different from image synthesis evaluation based on perceptual quality.
@@
Appendix C Additional Analysis of Latent OOD Properties
C.1 Background of Deterministic Diffusion
Our analytical studies and methodology designs are built upon a specific variant of diffusion formulations, i.e., the deterministic diffusion process. While the original diffusion denoising probabilistic models (DDPMs) involve a stochastic process for data generation via denoising (i.e., the same latent encoding will output different denoised images every time after the same generative chain), there is a variant of diffusion model that allows us to perform the denoising process in a deterministic way, known as the Denoising Diffusion Implicit Models (DDIMs) [song2020ddim]. DDIMs were initially proposed for the purpose of speeding up the denoising process, however, later research works extend DDIMs from faster sampling application to other usages including the inversion technique to convert a raw image to its arbitrary latent space in a deterministic and tractable way. As briefly stated in our main paper, the core theoretical difference between DDIMs and DDPMs lies within the nature of forward process, which modifies a Markovian process to a non-Markovian one.
@@ C.1 Background of Deterministic Diffusion
The choice of mean function from Eqn. 2 ensures that $q_{\sigma}(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\sqrt{\alpha_{t}}\mathbf{x}_{0},(1-\alpha_{t})\mathbf{I})$ for all $t$, so that it defines a joint inference distribution that matches the “marginals” as desired. The non-Markovian forward process can be derived from Bayes’ rule:
@@
In fact, in the original paper, the authors also explicitly stated that: “ The forward process from Eqn. 3 is also Gaussian (although we do not use this fact for the remainder of this paper)”.  [^1] This paper refer to the DDIM paper [song2020ddim]. While this Gaussian property was not emphasized and leveraged in the DDIMs paper, we find it useful in our context to explore the representation and generalization ability of pre-trained DDPMs.
In particular, the hyper-parameters for Gaussian scheduling $\alpha$ and $\beta$ in the context of DDIMs are slightly different from the original formulation in DDPMs [ho2020dpm]. Denote the original sequences from DDPMs as $\alpha^{\prime}_{t}$, then the $\alpha_{t}$ in this work follows the definition of DDIMs to be $\alpha_{t}=\prod^{T}_{t=1}\alpha^{\prime}_{t}$.
In addition to DDIMs, we note that the score-based formulation has also recently marked a deterministic variant, namely the Consistency Models [song2023consistency]. The core idea of the consistency model is, to some extent, similar to DDIMs, which allows the vanilla score-based stochastic diffusion models to achieve “one-step” denoising, by following the marginal distributions.
@@
Table 2: Reconstruction results for arbitrary images via deterministic diffusion. We use an iDDPM [nichol2021improveddmp] trained on AFHQ-Dog and 1K testing OOD images to compute the MAE (mean absolute error) reconstruction metric. Note DDIMs [song2020ddim] was initially proposed to accelerate DDPMs sampling, but have not been studied in this OOD reconstruction setting.
@@
pSp [richardson2021encoding] CelebA (ID) 0.079
e4e [tov2021designing] CelebA (ID) 0.092
ReStyle [alaluf2021restyle] CelebA (ID) 0.089
HFGI [wang2022high] CelebA (ID) 0.062
DDIMs [song2020ddim] Dog (ID) 0.073 $\pm$ 6e-4
@@
C.2 Latent Representation Ability via Reconstruction
In our work, we evaluate the latent representation capability through reconstruction experiments and report the corresponding quantitative results in Tab. 2. These experiments serve as a sanity check under the assumption that, given plausible noisy latent encodings, the model should at least be able to faithfully reconstruct arbitrary target OOD images.
C.3 Parameter-Independent Properties: Gaussian Priors
We seek a theoretically grounded explanation to the generalization properties of pre-trained DDPMs after the inversion. The takeaway message is: In theory, the inverted latent encodings also establish Gaussian priors as presented in Lemma 4.222However, in practice, due to the fact that pre-trained DMs themselves are function approximators, the samples after inversion do not establish perfect Gaussians but rather approximations.
Lemma C.1